### v1

# Open Heavy flavour mesons in hot asymmetric strange hadronic matter – a QCD sum rule approach

Abstract
The in-medium masses of the pseudoscalar open charm ($D$, $\bar{D}$, $D_{s}$ and $\bar{D_{s}}$) and open bottom ($B$, $\bar{B}$, $B_{s}$ and $\bar{B_{s}}$) mesons in hot asymmetric strange hadronic matter are studied within a QCD sum rule approach. These are computed using the medium modifications of the light quark condensates ($\langle{\bar{q_{i}}}{q_{i}}\rangle$, with $(q_{i},i=1,2,3)\equiv(u,d,s)$), the scalar gluon condensate, $\left\langle\frac{\alpha_{s}}{\pi}G^{a}_{\mu\nu}{G^{a}}^{\mu\nu}\right\rangle$, the twist-2 tensorial gluon operator, $\left\langle\frac{\alpha_{s}}{\pi}\Big{(}G^{a}_{\mu\sigma}{{G^{a}}_{\nu}}^{\sigma}u^{\mu}u^{\nu}-\frac{1}{4}G^{a}_{\mu\sigma}{G^{a}}^{\mu\sigma}\Big{)}\right\rangle$ (where $u^{\mu}$ is the 4-velocity of the medium) and other operators in the operator product expansion upto mass dimension 5. Within a chiral SU(3) model, the quark condensates are obtained from the medium modifications of the non-strange and strange scalar-isoscalar fields ($\sigma$ and $\zeta$), and the scalar-isovector field, $\delta$, whereas, the gluon (scalar and twist-2) condensates are computed from the in-medium value of the dilaton field, $\chi$, which is incorporated within the chiral model to mimic the broken scale invariance of QCD. The mixed quark-gluon condensate $\langle\bar{q_{i}}g_{s}\sigma.Gq_{i}\rangle$ is calculated from the in-medium quark condensate $\langle\bar{q_{i}}q_{i}\rangle$. The splittings of the masses of the $D-\bar{D}$ ($B-\bar{B})$, as well as $D_{s}-\bar{D_{s}}(B_{s}-\bar{B_{s}})$ in the hadronic medium are due to the odd part of the spectral function. For the strange-charm (strange-bottom) mesons ($(q_{i},i=3)\equiv s$), the value of $\langle{q_{i}}^{\dagger}{q_{i}}\rangle$ dominates over the ocntributions from the other operators, $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$ of the odd part of the sectral function. This is observed as opposite behaviour of the particle-antiparticle mass splittings with the mass of the particle to be larger (smaller) than the antiparticle mass in nuclear (hyperonic) matter. The effects of density, isospin asymmetry, strangeness and temperature on the masses of the open charm and open bottom mesons are observed to be appreciable.
I Introduction
The study of the properties of hadrons under extreme conditions of temperature and/or density is an important subject of research in strong interaction physics. The topic is of direct relevance for the experimental observables of the relativistic heavy-ion collision experiments, as well as, is relevant for study of the bulk properties of compact objects, e.g., the neutron stars. The light vector mesons have been extensively studied in the literature, as the medium modifications of these mesons can be observed in the dilepton spectra arising from heavy ion collision experiments Rapp_Wambach . In the recent years, a lot of studies on the in-medium properties of heavy flavour hadrons have been initiated due to its relevance and possible observability in the relativistic heavy ion collision experiments Prog_Nucl_Part_Phys_2017_Hosaka_et_al . The spectral modification of the charm hadrons will be investigated via $\bar{p}$-A collisions at FAIR at the future facility at GSI, Darmstadt Yamagata-Sekihara_et_al ; Rapp and charm production using pion beams on proton in J-PARC Kim1 ; Kim2 ; Garzon . The charm (bottom) meson-nucleon interaction can also lead to possibility of formation of charm (bottom) mesons bound in atomic-nuclei.
The properties of heavy flavour mesons in the (hot) hadronic matter have been studied extensively in the literature. These have been investigated using the potential models potential_models , the coupled channel approach coupled_channel_approach , the quark meson coupling (QMC) model QMC_models ; krein_jpsi ; N_Quarkonia_krein heavy quark symmetry and interaction of these mesons with nucleons via pion exchange Yasui_Sudoh_pion , heavy meson effective theory Yasui_Sudoh_heavy_meson_Eff_th , studying the heavy flavour meson as an impurity in nuclear matter Yasui_Sudoh_heavy_particle_impurity , and the QCD sum rule approach open_heavy_flavour_qsr_1 ; kimlee ; klingl ; amarvjpsi_qsr ; moritalee_1 ; moritalee_2 ; moritalee_3 ; moritalee_4 ; leemorita2 ; open_heavy_flavour_qsr_2 ; open_heavy_flavour_qsr_3 ; open_heavy_flavour_qsr_4 ; Wang_heavy_mesons_1 ; Wang_heavy_mesons_2 ; arvind_heavy_mesons_QSR_1 ; arvind_heavy_mesons_QSR_2 ; arvind_heavy_mesons_QSR_3 . Using leading order QCD formula pes1 ; pes2 ; voloshin , the medium modifications of the masses of the charmonium states due to the medium change of the scalar gluon condensate were studied in a linear density approximation in Ref.leeko , which showed much larger mass shifts for the excited states ($\psi(3686)$ and $\psi(3770)$), as compared to the mass shift of $J/\psi$. Within a chiral effective model sche1 ; heide1 ; paper3 ; kristof1 ; hartree , the in-medium masses of the charmonium (bottomonium) states amarvdmesonTprc ; amarvepja ; AM_DP_upsilon as well as of the open heavy flavor mesons amdmeson ; amarindamprc ; amarvdmesonTprc ; amarvepja ; DP_AM_Ds ; DP_AM_bbar ; DP_AM_Bs have been studied. Within the chiral effective model, the mass modifications of the quarkonium states arise due to the medium modification of a scalar dilaton field, which is introduced in the model to mimic the broken scale invariance of QCD. The open heavy flavour mesons have been studied due to their interactions with the baryons and the scalar mesons in the hadronic medium. The model has been used to study the partial decay widths of the heavy quarkonium states to the open heavy flavour mesons, in the (magnetized) hadronic medium amarvepja ; charmdecay_mag using a light quark pair creation model friman , namely the ${}^{3}P_{0}$ model 3p0_1 ; 3p0_2 ; 3p0_3 ; 3p0_4 as well as using a field theoretical model for composite hadrons amspmwg ; amspm_upsilon ; charmdw_mag ; upslndw_mag . Within the chiral effective model, in the presence of an external magnetic field, the effects of magnetized Dirac sea have been investigated on the properties of heavy flavour mesons in the hadronic medium HQ_DW_DS_AM_SPM_2023 ; Charmonium_DS ; Open_Bottom_DS The magnetized Dirac sea, accounting for the anomalous magnetic moment of nucleons is observed to lead to the phenomenon of inverse magnetic catalysis at finite densities for symmetric as well as asymmetric nuclear matter in presence of strong magnetic fields. Recently, the effects of the Dirac sea as well as mixing of the pseudoscalar and vector mesons on the masses, decay widths and the production cross-sections of the charmonium states in hot magnetized strange hadronic matter have been investigated in Ref. AMAKSPM24 .
Within the chiral SU(3) model, using the medium changes of the light quark condensates ($\langle\bar{u}u\rangle$, $\langle\bar{d}d\rangle$ and $\langle\bar{s}s\rangle$), obtained from the non-strange isoscalar ($\sigma$) isovector ($\delta$) and strange ($\zeta$) scalar fields and gluon condensates calculated from the scalar dilaton field, $\chi$, the light vector mesons ($\omega$, $\rho$ and $\phi$), in (magnetized) hadronic matter have been studied within the framework of QCD sum rule approach am_vecmeson_qsr ; vecqsr_mag . The effects of the magnetized Dirac sea on the spectral properties of light vector and axial-vector mesons in hot magnetized nuclear matter Spectr_rho_A1_DS and of the ground state heavy quarkonia QCDSR_Heavy_Quarkonia have also been studied with ths QCD sum rule approach. In the present work, we investigate the masses of the open charm ($D$, $\bar{D}$, $D_{s}$) and open bottom ($B$, $\bar{B}$, $B_{s}$) mesons in isospin asymmetric strange hadronic matter at finite temperature. In section II, we briefly describe the QCD sum rule approach used to compute the masses of the open charm and open bottom mesons, retaining the operators upto dimension 5 in the operator product expansion of the spectral function. Unlike the case of vacuum, in the hadronic medium, there is splitting between the masses of the particles and their antiparticles arising due to the odd part of the spectral function. In section III, we describe the chiral SU(3) model used to compute the in-medium values of the quark (light and strange), the gluon (scalar and twist-2) and the mixed quark-gluon operators, which are used for the calculation of the in-medium masses of the pseudoscalar open charm and open bottom mesons. Section IV discusses the results and in section V, we summarize the findings of the present work.
II QCD sum rule approach
The QCD sum rule approach relates the hadronic properties, encoded in spectral functions, to the QCD condensates Cohen_Prog_Part_Nucl_Phys . The basic ingredient is the two-point correlation function given as
$\Pi(q)=i\int d^{3}xe^{-iq.x}\langle\Omega|T(j(x)j(0)^{\dagger})|\Omega\rangle,$ (1)
where, $|\Omega\rangle$ is the physical ground state. Unlike the case of vacuum, the two-point correction function in the hadronic medium has contributions which are odd in $q_{0}$. This can be written in terms of even and odd functions of $q_{0}$ as
$\Pi(q_{0},\vec{q})=\Pi^{e}(q_{0},\vec{q})+q_{0}\Pi^{o}(q_{0},\vec{q}),$ (2)
with $\Pi^{e}(q_{0},\vec{q})=\frac{1}{2}(\Pi(q_{0},\vec{q})+\Pi(-q_{0},\vec{q}))$ and $\Pi^{o}(q_{0},\vec{q})=\frac{1}{2q_{0}}(\Pi(q_{0},\vec{q})-\Pi(-q_{0},\vec{q}))$. It might be noted here that $\Pi^{e(o)}(q_{0},\vec{q})$ depends on $q_{0}^{2}$. The Borel transformed sum rules for the even and odd parts are given as open_heavy_flavour_qsr_2
$\Pi^{e}_{OPE}(M^{2})=\Bigg{[}\int_{-\infty}^{s_{0}^{-}}+\int_{s_{0}^{-}}^{s_{0}^{+}}+\int_{s_{0}^{+}}^{\infty}\Bigg{]}\frac{s\;{\rm d}s\;{\rm Im}\;\Pi(s)}{\pi}e^{-{s}/{M^{2}}}$ (3)
and
$\Pi^{o}_{OPE}(M^{2})=\Bigg{[}\int_{-\infty}^{s_{0}^{-}}+\int_{s_{0}^{-}}^{s_{0}^{+}}+\int_{s_{0}^{+}}^{\infty}\Bigg{]}\frac{{\rm d}s\;{\rm Im}\;\Pi(s)}{\pi}e^{-{s}/{M^{2}}}.$ (4)
In the above equations, the subscript ‘$OPE$’ denotes the operator product exapnsion $\langle\Omega|T(j(x)j(0)^{\dagger})|\Omega\rangle=\sum_{n}C_{n}(x-y)\langle\Omega|O_{n}|\Omega\rangle$, with $\langle\Omega|O_{n}|\Omega\rangle$ as the QCD condensates for the operators $O_{n}$ and $C_{n}(x-y)$ are the Wilson coefficients, $s_{0}^{\pm}$ are the continuum thresholds, for the particle (antiparticle) and $M$ is the Borel mass. The open charm and open bottom mesons have been studied in symmetric nuclear matter within the QCD sum rue approach keeping operators upto dimension 5 in the OPE side of the spectral functions in Ref. open_heavy_flavour_qsr_2 . In the present work, we study the in-medium masses of these pseudoscalar mesons in hot isospin asymmetric strange hadronic matter. These spectral functions are obtained using the current operators for the particles ($p_{i}$), $j_{i}=j^{p}_{i}\equiv i{\bar{q_{i}}}\gamma_{5}Q$, with $(q_{i},i=1,2,3)\equiv(u,d,s)$, where, the heavy quark $Q=c$ corresponds to $D^{0}$, $D^{+}$ and $D_{s}^{+}$ mesons, and, $Q=b$ corresponds to $B^{-}$, $\bar{B^{0}}$ and $\bar{B_{s}}^{0}$ mesons respectively. The current operator is $j_{i}^{\dagger}\equiv i{\bar{Q}}\gamma_{5}{q_{i}}$, for the antiparticles ($\bar{D^{0}}$, $D^{-}$, $D_{s}^{-}$ for charm sector and $B^{+}$, $B^{0}$ and ${B_{s}}^{0}$ for the bottom sector). The Borel transformed even and odd parts of the OPE upto dimension 5 are obtained as open_heavy_flavour_qsr_2
$\Pi^{e}_{OPE}(M^{2})=\frac{1}{\pi}\int_{m_{Q}^{2}}^{\infty}{\rm Im}\Pi^{per}_{P}(s,\vec{q}=0)+e^{-{m_{Q}^{2}}/{M^{2}}}P^{e}(M^{2})$ (5)
with
$P^{e}(M^{2})$ $=$ $-m_{Q}\langle{\bar{q_{i}}}{q_{i}}\rangle+\frac{1}{2}m_{q_{i}}\langle{\bar{q_{i}}}{q_{i}}\rangle\Big{(}1+\frac{m_{Q}^{2}}{M^{2}}\Big{)}+\frac{m_{Q}}{2M^{2}}\Big{(}\frac{m_{Q}^{2}}{2M^{2}}-1\Big{)}\langle\bar{q_{i}}g_{s}\sigma\cdot Gq_{i}\rangle$ (6)
$+$ $\frac{1}{12}\langle\frac{\alpha_{s}}{\pi}G_{\mu\nu}G^{\mu\nu}\rangle+\Bigg{[}\Big{(}\frac{7}{18}+\frac{1}{3}\ln\Big{(}\frac{\mu^{2}m_{Q}^{2}}{M^{4}}\Big{)}-\frac{2\gamma_{E}}{3}\Big{)}\Big{(}\frac{m_{Q}^{2}}{M^{2}}-1\Big{)}-\frac{2}{3}\frac{m_{Q}^{2}}{M^{2}}\Bigg{]}$
$\times$ $\langle\frac{\alpha_{s}}{\pi}\Big{(}G^{a}_{\mu\sigma}{{G^{a}}_{\nu}}^{\sigma}u^{\mu}u^{\nu}-\frac{1}{4}G^{a}_{\mu\sigma}{G^{a}}^{\mu\sigma}\Big{)}\rangle+2\Big{(}\frac{m_{Q}^{2}}{M^{2}}-1\Big{)}\langle{q_{i}}^{\dagger}iD_{0}q_{i}\rangle$
$+$ $4\Big{(}\frac{m_{Q}^{3}}{2M^{4}}-\frac{m_{Q}}{M^{2}}\Big{)}\Big{[}\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle-\frac{1}{8}\langle\bar{q_{i}}g_{s}\sigma\cdot Gq_{i}\rangle\Big{]}$
and,
$\Pi^{o}_{OPE}(M^{2})=e^{-{m_{Q}^{2}}/{M^{2}}}P^{o}(M^{2})$ (7)
with
$P^{o}(M^{2})=\langle{q_{i}}^{\dagger}{q_{i}}\rangle-4\Big{(}\frac{m_{Q}^{2}}{2M^{4}}-\frac{1}{M^{2}}\Big{)}\langle{q_{i}}^{\dagger}D_{0}^{2}{q_{i}}\rangle-\frac{1}{M^{2}}\langle{q_{i}}^{\dagger}g_{s}\sigma\cdot Gq_{i}\rangle.$ (8)
In equation (6), the first and second terms arise due to the quark-antiquark condensate ($\langle{\bar{q_{i}}}q_{i}\rangle$), retaining the term proportional to the light quark mass ($m_{q_{i}}$) open_heavy_flavour_qsr_1 as given by the second term. The first term, due to the large mass of the heavy quark ($m_{Q}$) is the dominant contribution to the mass of the open charm (bottom) meson, arising from the dimension 3 quark-antiquark condensate in the operator product exapansion. The strong coupling constant, $\alpha_{s}$ is given as $\alpha_{s}=4\pi/\Big{[}\big{(}(11-2N_{f}/3)\rm{ln}(\mu^{2}/\Lambda_{QCD}^{2})\big{)}\Big{]}$, with $\Lambda_{QCD}=500$ MeV, and, $N_{f}$=4 (5) for the study of the open charm (bottom) mesons. $\mu$, the renormalization scale, is taken to be mass of the largest quark in the system, i.e, the mass of the charm (bottom) quark, for investigating the mass of the open charm (bottom) mesons open_heavy_flavour_qsr_2 , and $\gamma_{E}$ is the Euler constant. The other terms are the higher mass dimension operators, which include the dimension-4 scalar and twist-2 gluon condensate, the mixed quark-gluon operator (the abbreviation $\sigma\cdot G$ is used for $\sigma_{\mu\nu}G^{\mu\nu}$), and terms proportional to the operators $\langle{q_{i}}^{\dagger}iD_{0}q_{i}\rangle$ and $\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle$. In the odd part of the Borel transformed spectral function, the dominant contribution in eqn (8) is from the number density of the quark ($\langle{q_{i}}^{\dagger}{q_{i}}\rangle$), along with the contributions from the dimension 5 operators, $\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma\cdot Gq_{i}\rangle$. In the term proportional to the twist-2 gluon condensate operator, $u^{\mu}$ is the 4-velocity of the hadronic medium. For the spectral function, the ansatz
${\rm Im}\Pi(s)=\pi F_{+}\delta(s-m_{+})-\pi F_{-}\delta(s+m_{-}),$ (9)
is used, where, $m_{\pm}$ correspond to the mass of the particle (antiparticle). We choose the continuum thresholds as $s_{0}^{\pm}=\pm s_{0}$, similar to the case of vacuum. The masses for the particles and antiparticle, $m_{\pm}$, corresponding to the open heavy flavour mesons, are solved by using the spectral function given by equation (9) in the equations (3) and (4) open_heavy_flavour_qsr_2 . The expressions for the masses of the particle and antiparticle are given as where,
$m_{\pm}=\Bigg{[}\frac{1}{4}\Big{(}\frac{gf^{\prime}-fg^{\prime}}{f^{2}+gg^{\prime}}\Big{)}^{2}-\Big{(}\frac{ff^{\prime}+{g^{\prime}}^{2}}{f^{2}+gg^{\prime}}\Big{)}\Bigg{]}^{1/2}\pm\frac{1}{2}\Big{(}\frac{gf^{\prime}-fg^{\prime}}{f^{2}+gg^{\prime}}\Big{)}$ (10)
where, $f=\Pi_{OPE}^{e}(M^{2})$ and $g=\Pi_{OPE}^{o}(M^{2})$, the even and odd parts of the spectral function, are given by equations (5) and (7), and, $f^{\prime}$ and $g^{\prime}$ denote the derivatives of $f$ and $g$ with respect to $1/{M^{2}}$. As can be observed from eqn (10), the mass difference between the particles ($D^{0}$, $D^{+}$ and $D_{s}^{+}$) and their antiparticles ($\bar{D^{0}}$, ${D^{-}}$ and $D_{s}^{-}$) for the charm sector, and between the particles ($B^{-}$, $\bar{B^{0}}$ and ${\bar{B_{s}}}^{0}$) and their antiparticles ($B^{+}$, $B^{0}$ and $B_{s}^{0}$) for the bottom sector, in the hadronic medium arise from the values of $\langle{q_{i}}^{\dagger}{q_{i}}\rangle$, $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$ of the odd part of the spectral function ($g$), hence due to the expectation values of the QCD operators $\langle{q_{i}}^{\dagger}{q_{i}}\rangle$, $\langle{q_{i}}^{\dagger}D_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma\cdot Gq_{i}\rangle$ in $P^{o}(M^{2})$ given by equation (8). The in-medium QCD condensates are calculated within a chiral SU(3) model, as described in the following section.
III Chiral SU(3) model
In this section, we briefly describe the chiral SU(3) model paper3 ; hartree ; kristof1 ; AMAKSPM24 which is used to obtain the in-medium QCD condenstaes in the hot asymmetric strange hadronic matter. The expectation values of these QCD operators are central to the calculation of the in-medium masses of the open charm and open bottom mesons within the framework of the QCD sum rule approach, as described in the previous section. The chiral SU(3) model paper3 ; hartree ; kristof1 is based on a non-linear realization of chiral symmetry Weinberg ; coleman ; Bardeen and incorporates the broken scale invariance of QCD through introduction of a scalar dilaton field, $\chi$ sche1 ; heide1 . The Lagrangian density of chiral SU(3) model is written as paper3
${\cal L}={\cal L}_{kin}+\sum_{W=X,Y,V,A,u}{\cal L}_{BW}+{\cal L}_{vec}+{\cal L}_{0}+{\cal L}_{\chi}+{\cal L}_{SB}.$ (11)
In Eq.(11), ${\cal L}_{kin}$ is the kinetic energy term, ${\cal L}_{BW}$ is the baryon-meson interaction term, where, $W=X,Y,V,A,u$ correspond to the interactions of the baryons with the scalar, pseudoscalar singlet, vector meson, axial-vector meson and pseudoscalar octet mesons respectively. The baryons-spin-0 meson interaction term generates the baryon masses. ${\cal L}_{vec}$ describes the dynamical mass generation of the vector mesons via couplings to the scalar mesons and contain additionally quartic self-interactions of the vector fields. ${\cal L}_{0}$ contains the meson-meson interaction terms, ${\cal L}_{\chi}$ is the QCD scale breaking term in the Lagrangian density, including a logarthimic potential of the scalar dilaton field, $\chi$, and, ${\cal L}_{SB}$ describes the explicit chiral symmetry breaking. In the present study of isospin asymmetric strange hadronic matter at finite temperature, the mean field approximation is used, in which the meson fields are replaced by their expectation values. The meson fields which have non-zero expectation values are the scalar ($\sigma,\zeta,\delta)$ and the vector fields, $\omega^{\mu}\rightarrow\omega\delta^{\mu 0}$, $\rho^{\mu a}\rightarrow\rho\delta^{\mu 0}\delta_{a3}$ and, $\phi^{\mu}\rightarrow\phi\delta^{\mu 0}$, and, the expectation values of the other mesons are zero. We also use the approximations $\bar{\psi}_{i}\psi_{j}\rightarrow\delta_{ij}\langle\bar{\psi}_{i}\psi_{i}\rangle\equiv\delta_{ij}\rho_{i}^{s}$ and $\bar{\psi}_{i}\gamma^{\mu}\psi_{j}\rightarrow\delta_{ij}\delta^{\mu 0}\langle\bar{\psi}_{i}\gamma^{0}\psi_{i}\rangle\equiv\delta_{ij}\delta^{\mu 0}\rho_{i}$, where, $\rho_{i}^{s}$ and $\rho_{i}$ are the scalar and number density of baryon of species, $i$. In the above approximation, the terms which contribute to the baryon-meson interaction, ${\cal L}_{BW}$ of equation (11) arise from the interactions due to the scalar ($S$) and the vector ($V$) mesons, and, is given as
${\cal L}_{BS}+{\cal L}_{BV}=\sum_{i}{\bar{\psi}}^{i}(g_{\sigma i}\sigma+g_{\zeta i}\zeta+g_{\delta i}\delta-g_{\omega i}\gamma^{0}\omega-g_{\rho i}\gamma^{0}\rho-g_{\phi i}\gamma^{0}\phi)\psi^{i}.$ (12)
The explicit symmetry breaking term in the chiral model is given by
${\cal L}_{SB}$ $=$ $-\left(\frac{\chi}{\chi_{0}}\right)^{2}{\rm Tr}\left[{\rm diag}\left(\frac{m_{\pi}^{2}f_{\pi}}{2}(\sigma+\delta),\frac{m_{\pi}^{2}f_{\pi}}{2}(\sigma-\delta),\Big{(}\sqrt{2}m_{K}^{2}f_{K}-\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\Big{)}\zeta\right)\right],$ (13)
which reduces to eqn. (8), and in QCD, this term is given as ${\cal L}_{SB}^{QCD}=-Tr[{\rm diag}(m_{u}\bar{u}u,m_{d}\bar{d}d,m_{s}\bar{s}s)]$. In the above, the matrices, whose traces give the corresponding Lagrangian densities have been explicitly written down, and a comparison of their matrix elements relates the quark condensates to the values of the scalar fields as am_vecmeson_qsr
$m_{u}\langle\bar{u}u\rangle$ $=$ $\Big{(}\frac{\chi}{\chi_{0}}\Big{)}^{2}\frac{m_{\pi}^{2}f_{\pi}}{2}(\sigma+\delta),\;m_{d}\langle\bar{d}d\rangle=\left(\frac{\chi}{\chi_{0}}\right)^{2}\frac{m_{\pi}^{2}f_{\pi}}{2}(\sigma-\delta),$
$m_{s}\langle\bar{s}s\rangle$ $=$ $\left(\frac{\chi}{\chi_{0}}\right)^{2}\big{(}\sqrt{2}m_{K}^{2}f_{K}-\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\big{)}\zeta.$ (14)
The medium modifications of the quark condensates $\langle\bar{q_{i}}q_{i}\rangle$, with $q_{i}\equiv u,d,s$ for $i=1,2,3$, are obtained from the medium modifications of the scalar fields ($\sigma$ and $\zeta$ and $\delta$).
Figure 1: Masses of $D^{\pm}$ (in panels (a), (c) and (e)) and $D^{0}(\bar{D^{0}})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
Figure 2: Same as Fig. 1, with $f_{s}$=0.3.
Figure 3: Mass shifts of $D^{\pm}$ (in panels (a), (c) and (e)) and $D^{0}(\bar{D^{0}})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the baryon density in units of nuclear saturation density for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines).
Figure 4: Same as fig. 3, with $f_{s}$=0.3.
Figure 5: Masses of $D_{s}^{\pm}$ mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV in panels (a), (c) and (e) and in panels (b), (d) and (f), in nuclear ($f_{s}$=0) and hyperonic (with $f_{s}$=0.3) matter for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
Figure 6: Mass shifts of $D_{s}^{\pm}$ mesons are plotted as functions of the baryon density in units of nuclear saturation density for $\eta=0.5$ (in panels (a), (c) and (e)) for nuclear matter ($f_{s}$=0) and in panels (b), (d) and (f) for strange hadronic matter with $f_{s}$=0.3) for T=50, 100 and 150 MeV and compared to the case of $\eta$=0 (shown as dotted lines).
Figure 7: Masses of $B^{+}({B^{-}})$ (in panels (a), (c) and (e)) and $B^{0}(\bar{B}^{0})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
Figure 8: Same as fig. 7, with $f_{s}$=0.3.
Figure 9: Mass shifts of $B^{+}({B^{-}})$ (in panels (a), (c) and (e)) and $B^{0}(\bar{B}^{0})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the baryon density in units of nuclear saturation density for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines).
Figure 10: Same as fig. 9, with $f_{s}$=0.3.
Figure 11: Masses of $B_{s}^{0}$ and ${\bar{B}}_{s}^{0}$ mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV for nuclear matter ($f_{s}$=0) in panels (a), (c) and (e) and for hyperonic (with $f_{s}$=0.3) in panels (b), (d) and (f), for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
Figure 12: Mass shifts of $B_{s}^{0}$ and ${\bar{B}}_{s}^{0}$ mesons are plotted as functions of the baryon density in units of nuclear saturation density for $\eta=0.5$ (in panels (a), (c) and (e)) for nuclear matter ($f_{s}$=0) and in panels (b), (d) and (f) for strange hadronic matter with $f_{s}$=0.3) for T=50, 100 and 150 MeV and compared to the case of $\eta$=0 (shown as dotted lines).
The scalar gluon condensate ($\left\langle\frac{\alpha_{s}}{\pi}G^{a}_{\mu\nu}{G^{a}}^{\mu\nu}\right\rangle$ and the expectation value of the twist-2 tensorial gluon operator, $\left\langle\frac{\alpha_{s}}{\pi}\Big{(}G^{a}_{\mu\sigma}{{G^{a}}_{\nu}}^{\sigma}u^{\mu}u^{\nu}-\frac{1}{4}G^{a}_{\mu\sigma}{G^{a}}^{\mu\sigma}\Big{)}\right\rangle$, are obtained from the in-medium value of the dilaton field, $\chi$ amarvjpsi_qsr , which is incorporated in the chiral SU(3) model to simulate the broken scale invariance of QCD. The Lagrangian density for the scalar dilaton field, $\chi$, is given as amarvjpsi_qsr
${\cal L}_{\chi}=\frac{1}{2}(\partial_{\mu}\chi)(\partial^{\mu}\chi)-k_{4}\chi^{4}-\frac{1}{4}\chi^{4}{\rm{ln}}\Bigg{(}\frac{\chi^{4}}{\chi_{0}^{4}}\Bigg{)}+\frac{d}{3}\chi^{4}{\rm{ln}}\Bigg{(}\bigg{(}\frac{\left(\sigma^{2}-\delta^{2}\right)\zeta}{\sigma_{0}^{2}\zeta_{0}}\bigg{)}\bigg{(}\frac{\chi}{\chi_{0}}\bigg{)}^{3}\Bigg{)},$ (15)
where, the last two terms in the above Lagrangian density are the scale breaking terms. The energy momentum tensor obtained from the above Langrangian density is given as
$T_{\mu\nu}=(\partial_{\mu}\chi)\Bigg{(}\frac{\partial{{\cal L}_{\chi}}}{\partial(\partial^{\nu}\chi)}\Bigg{)}-g_{\mu\nu}{\cal L}_{\chi}\equiv(\partial_{\mu}\chi)(\partial_{\nu}\chi)-g_{\mu\nu}{\cal L}_{\chi}.$ (16)
Accounting for the finite mass of the quarks, the energy momentum of the QCD can be written in terms of a symmetric traceless part and a trace part as
$T_{\mu\nu}=-ST({G^{a}}_{\mu\sigma}{{G^{a}}_{\nu}}^{\sigma})+\frac{g_{\mu\nu}}{4}\Big{(}\frac{\beta_{QCD}}{2g}G_{\sigma\kappa}^{a}{G^{a}}^{\sigma\kappa}+\sum_{i}m_{i}\bar{q_{i}}q_{i}\Big{)},$ (17)
Writing
$\langle\frac{\alpha_{s}}{\pi}{G^{a}}_{\mu\sigma}{{G^{a}}_{\nu}}^{\sigma}\rangle=\Big{(}u_{\mu}u_{\nu}-\frac{g_{\mu\nu}}{4}\Big{)}G_{2},$ (18)
where $u_{\mu}$ is the 4-velocity of the nuclear medium, taken as $u_{\mu}=(1,0,0,0)$, we obtain the energy momentum tensor in QCD as
$T_{\mu\nu}=-\Big{(}\frac{\pi}{\alpha_{s}}\Big{)}\Big{(}u_{\mu}u_{\nu}-\frac{g_{\mu\nu}}{4}\Big{)}G_{2}+\frac{g_{\mu\nu}}{4}\Big{(}\frac{\beta_{QCD}}{2g}{G^{a}}_{\sigma\kappa}{G^{a}}^{\sigma\kappa}+\sum_{i}m_{i}\bar{q_{i}}q_{i}\Big{)}.$ (19)
We calculate the expectation values of the scalar and twist-2 gluon operators by equating the energy momentum tensor for the dilaton field with that of QCD, multiplying both sides by $g^{\mu\nu}$ and $(u^{\mu}u^{\nu}-\frac{g^{\mu\nu}}{4})$ ($u^{\mu}=(1,0,0,0)$ is the 4-velocity of the hadronic medium at rest) respectively, and summing over the indices ($\mu$ and $\nu$) amarvjpsi_qsr . We then obtain amarvjpsi_qsr
${T_{\mu}}^{\mu}=\sum_{i}m_{i}\bar{q_{i}}q_{i}+\langle\frac{\beta_{QCD}}{2g}G_{\mu\nu}^{a}G^{\mu\nu a}\rangle=-(1-d)\chi^{4}$ (20)
and,
$\langle\frac{\alpha_{s}}{\pi}{G^{a}}_{\mu\sigma}{{G^{a}}_{\nu}}^{\sigma}\rangle(u^{\mu}u^{\nu}-\frac{g^{\mu\nu}}{4})\equiv\frac{3}{4}G_{2}=\frac{3\alpha_{s}}{4\pi}\Bigg{[}-(1-d+4k_{4})(\chi^{4}-{\chi_{0}}^{4})$ (21)
-- $\chi^{4}{\rm{ln}}\Big{(}\frac{\chi^{4}}{{\chi_{0}}^{4}}\Big{)}+\frac{4}{3}d\chi^{4}{\rm{ln}}\Bigg{(}\bigg{(}\frac{\left(\sigma^{2}-\delta^{2}\right)\zeta}{\sigma_{0}^{2}\zeta_{0}}\bigg{)}\bigg{(}\frac{\chi}{\chi_{0}}\bigg{)}^{3}\Bigg{)}\Bigg{]}.$
In equation (20), the first term of the trace of the energy-momentum tensor within the chiral SU(3) model is the negative of the explicit chiral symmetry breaking term, ${\cal L}_{SB}$ given by equation (13). It might be noted here that in deriving the relation of the twist-2 gluon operator with the dilaton field (given by eqn (21)),we have assumed that the dilaton field, $\chi$ is non-relativistic, and hence have assumed that $\langle(\partial_{i}\chi)(\partial_{i}\chi)\rangle_{medium}\simeq\langle(\partial_{i}\chi)(\partial_{i}\chi)\rangle_{vac}$. Using the QCD $\beta$ function at one loop level, for $N_{c}$ colors and $N_{f}$ flavors
$\beta_{\rm{QCD}}\left(g\right)=-\frac{11N_{c}g^{3}}{48\pi^{2}}\left(1-\frac{2N_{f}}{11N_{c}}\right),$ (22)
for $N_{c}$=3 and $N_{f}$=3, the scalar gluon condensate is obtained as
$\left\langle\frac{\alpha_{s}}{\pi}{G^{a}}_{\mu\nu}{G^{a}}^{\mu\nu}\right\rangle=\frac{8}{9}\Bigg{[}(1-d)\chi^{4}+\left(\frac{\chi}{\chi_{0}}\right)^{2}\left(m_{\pi}^{2}f_{\pi}\sigma+\big{(}\sqrt{2}m_{k}^{2}f_{k}-\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\big{)}\zeta\right)\Bigg{]}.$ (23)
The parameter $d$ in equation (20) originates from the second logarithmic term of equation (15). The parameter $d$, along with the other parameters corresponding to the meson-meson interaction Lagrangian density, ${\cal L}_{0}$ of the general Lagrangian given by (11), are fitted so as to ensure extrema in the vacuum for the $\sigma$, $\zeta$ and $\chi$ field equations, to reproduce the vacuum masses of the $\eta$ and $\eta^{\prime}$ mesons, the mass of the $\sigma$ meson around 500 MeV, and, pressure, p($\rho_{0}$)=0, with $\rho_{0}$ as the nuclear matter saturation density paper3 ; amarindamprc . The scalar fields (the non-strange isoscalar field, $\sigma$, the strange isoscalar field, $\zeta$ and the isovector field $\delta$), the dilaton field, $\chi$ and the vector fields (non-strange isoscalar $\omega$, strange isoscalar $\phi$ and non-strange isovector $\rho$) are obtained by minimizing the thermodynamic potential $\Omega$ for given values of temperature, baryonic density, $\rho_{B}$, the isospin asymmetry parameter, $\eta=-\frac{\Sigma_{i}I_{3i}\rho_{i}}{\rho_{B}}$, the strangeness fraction, $f_{s}=\frac{\Sigma_{i}|s_{i}|\rho_{i}}{\rho_{B}}$, where, $I_{3i}$ and $|s_{i}|$ denote the 3rd component of isospin quantum number and number of strange quarks in $i^{th}$ baryon, respectively.
The effects of the scalar and twist-2 gluon condensates, as obtained from the medium modification of the dilaton field, $\chi$, within the chiral SU(3) model, on the masses of the vector and pseudoscalar charmonium ground states ($J/\psi$ and $\eta_{c}$) were studied in Ref. amarvjpsi_qsr and on the masses of the heavy quarkonium (charmonium and bottomonium) states in Ref. QCDSR_Heavy_Quarkonia , using QCD sum rule approach. In the present work, the effects of the medium modifications of the QCD operators, upto dimension 5, on the masses of the open charm and open bottom mesons are investigated within the QCD sum rule approach. The non-strange ($\langle\bar{u}u\rangle$, $\langle\bar{d}d\rangle$) and strange ($\langle\bar{s}s\rangle$) quark condensates are obtained from the scalar fields using equation (14) and the scalar and twist-2 gluon condensates are calculated from the dilaton field, $\chi$, using equations (20 and (21) within the chiral SU(3) model. The dimension 5 mixed quark-gluon condensate $\langle\bar{q_{i}}g_{s}\sigma.G{q_{i}}\rangle$ in the operator product expansion is obtained from the in-medium quark condensate $\langle\bar{q_{i}}{q_{i}}\rangle$, with $i=1,2,3\equiv u,d,s$, as computed within the chiral SU(3) model, by using the relation $\langle\bar{q_{i}}g_{s}\sigma.G{q_{i}}\rangle=2\lambda_{q}^{2}\langle\bar{q_{i}}{q_{i}}\rangle$. The estimates for $\lambda_{q}^{2}$ Cohen_Prog_Part_Nucl_Phys in vacuum using standard QCD sum rule approach is 0.4 GeV2 Belyaev , the value estimated using lattice QCD is 0.55 GeV2 Kremer , and, $\lambda_{q}^{2}\sim$ 1.2 GeV2, in a QCD sum rule analysis of the pion wave function using non-local condensates Mikhailov as well as within an instanton liquid model Shuryak . Using the leading order approximations for the parton distribution functions Glueck_Reya_Vogt , the values of the higher order operators (or combinations of these operators) used in the present investigation are as follows Jin_Cohen_Furnstahl_Griegel_PRC47_2882_1983 . The estimates for $q_{i}=u,d$ are
$\langle q_{i}^{\dagger}iD_{0}q_{i}\rangle=0.18\;{\rm GeV}\rho_{B},$ (24)
$-\langle{\bar{q}_{i}}D_{0}^{2}q_{i}\rangle+\frac{1}{8}\langle\bar{q_{i}}g_{s}\sigma.G{q_{i}}\rangle=0.3\;{\rm GeV}^{2}\;\rho_{B},$ (25)
and
$-\langle q_{i}^{\dagger}D_{0}^{2}q_{i}\rangle+\frac{1}{12}\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle=0.031\;{\rm GeV}^{2}\;\rho_{B},$ (26)
for the case of symmetric nuclear matter. For the strange quark ($q_{i}=s$), the estimate used in the present investigation for the above operator is taken to be $\langle s^{\dagger}iD_{0}s\rangle=0.06\;{\rm GeV}\rho_{B}$ Cohen_Prog_Part_Nucl_Phys . For the consideration of the strange-heavy flavor mesons, the estimates of the combinations of the operators given by equations (25) and (26) (for $q_{i}$=s) are given as open_heavy_flavour_qsr_2
$-\langle{\bar{s}}D_{0}^{2}s\rangle+\frac{1}{8}\langle\bar{s}g_{s}\sigma.G{s}\rangle=y\Big{(}-\langle q_{i}^{\dagger}D_{0}^{2}q_{i}\rangle+\frac{1}{12}\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle\Big{)}$ (27)
and
$-\langle s^{\dagger}D_{0}^{2}s\rangle+\frac{1}{12}\langle{s}^{\dagger}g_{s}\sigma.G{s}\rangle=y\Big{(}-\langle q_{i}^{\dagger}D_{0}^{2}q_{i}\rangle+\frac{1}{12}\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle\Big{)},$ (28)
where, $q_{i}=u\;\;{\rm{or}}\;\;d$ in the RHS of the above equations, which have equal values for the operators for symmetric nuclear matter. In equations (27) and (28), $y$ is the strangeness content of the nucleon, given as
$y=\frac{\langle\bar{s}s\rangle_{med}}{\langle\bar{q}q\rangle_{med}},$ (29)
with $\langle\bar{s}s\rangle_{med}=\langle\bar{s}s\rangle-\langle\bar{s}s\rangle_{vac}$, and, ${\langle\bar{q}q\rangle}_{med}=({\langle\bar{u}u\rangle}_{med}+{\langle\bar{d}d\rangle}_{med})/2$ is the average of the quark-antiquark condensates for $u$ and $d$ quarks after subtracting the vacuum values. For the isospin asymmetric nuclear (hyperonic) matter as considered in the present investigation, we generalize the above relations as follows. For the non-strange quarks, $q_{i}=u,d$, for $i=1,2$,
$-\langle{\bar{q}_{i}}D_{0}^{2}q\rangle+\frac{1}{8}\langle\bar{q_{i}}g_{s}\sigma.G{q_{i}}\rangle=0.3\;{\rm GeV}^{2}\;\rho_{B}\cdot\frac{\langle\bar{q_{i}}q_{i}\rangle_{med}}{\langle\bar{q}q\rangle_{med}},$ (30)
$-\langle q_{i}^{\dagger}D_{0}^{2}q_{i}\rangle+\frac{1}{12}\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle=0.031\;{\rm GeV}^{2}\;\rho_{B}\cdot\frac{\langle\bar{q_{i}}q_{i}\rangle_{med}}{\langle\bar{q}q\rangle_{med}},$ (31)
For the QCD operators involving the strange quark, we take
$-\langle{\bar{s}}D_{0}^{2}s\rangle+\frac{1}{8}\langle\bar{s}g_{s}\sigma.G{s}\rangle=0.3\;{\rm GeV}^{2}\;\rho_{B}\cdot y,$ (32)
$-\langle s^{\dagger}D_{0}^{2}s\rangle+\frac{1}{12}\langle{s}^{\dagger}g_{s}\sigma.G{s}\rangle=0.031\;{\rm GeV}^{2}\;\rho_{B}\cdot y,$ (33)
with $y$, the strangeness content of the nucleon defined by equation (29). The estimate for the operator $\langle q_{i}^{\dagger}g_{s}\sigma.Gq_{i}\rangle=0.33\;{\rm{GeV}}^{2}\rho_{B}$, for $q_{i}=u\;\;{\rm{or}}\;\;d$ for the symmetric nuclear matter open_heavy_flavour_qsr_2 is generalized to the isospin asymmetric case in the present investigation as
$\langle q_{i}^{\dagger}g_{s}\sigma.Gq_{i}\rangle=0.33\;{\rm{GeV}}^{2}\rho_{B}\cdot\frac{\langle\bar{q_{i}}q_{i}\rangle_{med}}{\langle\bar{q}q\rangle_{med}},$ (34)
for $q_{i}=u\;\;{\rm{and}}\;\;d$, and, for the operator with strange quark given as,
$\langle{s}^{\dagger}g_{s}\sigma.G{s}\rangle=0.33\;{\rm GeV}^{2}\;\rho_{B}\cdot y.$ (35)
Using the estimates for the combinations of the operators given by equations 31 and 33, one obtaines the estimates for $\langle{q_{i}}^{\dagger}D_{0}^{2}{q_{i}}\rangle$, for $(q_{i},i=1,2,3)\equiv(u,d,s)$. Using the values for the QCD operators for the even and odd parts of the spectral function as obtained above, the masses of the open charm and open bottom mesons are computed using the QCD sum rule approach.
IV Results and discussions
In this section, we discuss the results obtained for the in-medium masses of the open charm ($D$, $\bar{D}$, $D_{s}$ and $\bar{D_{s}}$) and open bottom ($B$, $\bar{B}$, $B_{s}$ and $\bar{B_{s}}$) mesons, and, their antiparticles in asymmetric strange hadronic matter at finite temperatures. These are computed using the QCD sum rule approach retaining QCD operators upto dimension 5, with the medium dependence, e.g., the temperature, isospin asymmetry, strangeness and density dependence of these operators calculated using a chiral SU(3) model. In the present investigation, we choose the parameters as $m_{u}=m_{d}$=6 MeV, $m_{s}$=185 MeV, $m_{c}$=1.5 GeV, $m_{b}$=4.7 GeV and $\lambda_{q}$=1.2 GeV2). For the charm sector, the vacuum masses for the $D$ and $\bar{D}$ mesons are obtained to be 1869 MeV for $s_{0}$=7.5 GeV2, and, the mass of $D_{s}^{\pm}$ is obtained to be around 1916 MeV for $s_{0}$=15 GeV2. For the bottom sector, the vacuum masses for the $B$ and $\bar{B}$ mesons are obtained to be 5277.8 MeV for $s_{0}$=37.8 GeV2, and, the mass of $B_{s}^{0}({\bar{B}}_{s}^{0})$ is obtained to be around 5366.455 MeV for $s_{0}=42.3\;{\rm{GeV}}^{2}$. There is a splitting between the masses of the open heavy flavour particles ($D^{0}$, $D^{+}$, $D_{s}^{+}$, $B^{-}$, $\bar{B^{0}}$ and ${\bar{B_{s}}}^{0}$) and their antiparticles ($\bar{D^{0}}$, ${D^{-}}$, $D_{s}^{-}$, $B^{+}$, $B^{0}$ and $B_{s}^{0}$), which arises from the contributions from the number density of the $i$-th quark ($\langle{q_{i}}^{\dagger}{q_{i}}\rangle$), along with the contributions from the dimension 5 operators, $\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma\cdot Gq_{i}\rangle$ of the odd part of the spectral function.
In figures 1 and 2, we plot the masses of the charged and neutral open charm mesons as functions of the Borel mass, $M$, for the cases of nuclear matter ($f_{s}$=0) and strange hadronic matter (with $f_{s}$=0.3) for $\rho_{B}=2\rho_{0}$. These are shown for the isospin asymmetry parameter, $\eta$=0.5, and are compared to the case of isopsin symmetric case ($\eta$=0). The vacuum case is shown as the dotted line. At finite densities, plotted for $\rho_{B}=2\rho_{0}$ in these figures, the mass of the antiparticle $D^{-}(\bar{D^{0}}$) is observed to be higher than $D^{+}(D^{0})$ both in the hot nuclear matter, as well as in strange hadronic matter. As can be observed from these figures, the effect of isospin asymmetry is to reduce (enhance) the mass of $D^{+}(D^{0})$ within the iso-doublet ($D^{0}$, $D^{+}$), and, of ${\bar{D}}^{0}(D^{-})$ within the iso-doublet ($D^{-}$,${\bar{D}}^{0}$) for both nuclear matter as well as strange hadronic matter. The effects from isospin asymmetry are observed to be less to the particle $D^{0}$ as compared to its antiparticle $\bar{D^{0}}$.
The shifts in the masses of the charged ($D^{\pm}$) and neutral ($D^{0},{\bar{D}}^{0}$) from the vacuum values are plotted as functions of the baryon density in units of nuclear matter saturation density, in figures 3 and 4 for nuclear matter ($f_{s}$=0) and strange hadronic matter (with $f_{s}$=0.3) respectively. The effect from isospin asymmetry is observed to be appreciable for higher values of the baryon density, especially for $\bar{D}^{0}$ meson. In isospin symmetric ($\eta$=0) nuclear (hyperonic) matter, the masses of open charm mesons within the $D(D^{0},D^{+})$ and $\bar{D}(D^{-},\bar{D^{0}})$ doublets are degenerate. At $\rho_{B}=\rho_{0}$, the difference in the mass of $D^{+}(D^{0})$ ($\sim$ 1912 MeV) and the mass of $\bar{D^{0}}(D^{-})$ ($\sim$ 1944 MeV) is around 32 MeV, which may be compared to the mass difference of around 30 MeV in Ref. open_heavy_flavour_qsr_2 for the isospin symmetric ($\eta$=0) nuclear matter. In the presence of strangeness in the medium (with $f_{s}$=0.3), as plotted in figure 4, for isospin symmetric matter ($\eta$=0), the mass difference of $D^{0}(D^{+})$ ($\sim$ 1914.4 MeV) and $\bar{D^{0}}(D^{-})$ ($\sim$ 1935.9 MeV) is observed to be modified to 21.5 MeV. In the presence of isospin asymmetry, the mass splitting between the neutral (charged) particle and antiparticle, $D^{+}(D^{0})$ and $D-(\bar{D^{0}})$ is observed to increase (drop) with increase (drop) as the density is raised. However, the effect from the ispspin asymmetry is lessened in the presence of strangeness in the medium, as can be seen from figures 3 and 4.
In fgure 5, the masses of the strange-charm mesons, $D^{\pm}$ are plotted as functions of the Borel mass, $M$ for the nuclear and strange hadronic matter for $\rho_{B}=2\rho_{0}$. The Borel curve for these masses are shown for the vacuum (shown as a dot-dashed line). In nuclear matter, unlike the observation of the mass of the particle, $D^{+}(D^{0})$ to be smaller than the mass of the antiparticle, $D^{-}({\bar{D}}^{0})$ (as seen in figures 1 and 3, the opposite trend of the mass of particle ($D_{s}^{+}$) to be higher than the mass of $D_{s}^{+}$ is observed in nuclear matter. On the other hand, the non-zero value of $\langle s^{\dagger}s\rangle$ in the strange hadronic matter leads to the antiparticle ($D_{s}^{-}$) mass to be higher than that of $D_{s}^{+}$ in hyperonic medium. Note that $\langle q_{i}^{\dagger}q_{i}\rangle$ for the $D(\bar{D})$ mesons corresponds to the number density of the non-strange quarks ($u$ or $d$), whereas, for $D_{s}^{\pm}$, $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$, the number density of the $s$-quark, the latter is zero for nuclear matter. The mass splitting between the particle and antiparticle is thus mainly driven by the term $\langle q_{i}^{\dagger}q_{i}\rangle$, which dominates over the contributions from the dimension 5 operators $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$. In figure 6, the mass shifts of the strange-charm mesons are plotted as a function of the baryon density in units of nuclear matter saturation density. With isospin asymmetry, the mass shifts of both the $D_{s}^{\pm}$ are observed to be smaller as compared to the isospin symmetric case for nuclear matter. The trend is opposite in the presence of strangeness in the medium. The effect due to isospin asymmetry is observed to be marginal in strange hadronic matter for T=50 and 100 MeV, whereas, at the higher temperature, T=150 MeV, the effect is much larger.
The masses of the $B$ and $\bar{B}$ mesons are plotted as functions of the Borel mass in figures 7 and 8 for $\rho_{B}=2\rho_{0}$ for nuclear matter ($f_{s}$=0) and strange hadronic matter (with $f_{s}$=0.3) respectively. The density dependence of the mass shifts of $B$ and $\bar{B}$ mesons are shown in figures 9 and 10 for $f_{s}$=0 and $f_{s}$=0.3 respectively. Similar to the case of $D$ and $\bar{D}$ mesons (see figures 1,2, 3 and 4), the masses of the antiparticles are observed to be larger than the masses of the particles in nuclear matter as well as strange hadronic matter. For isospin symmetric nuclear (hyperonic) matter, the masses of the mesons within the $B$ and $\bar{B}$ doublets are degenerate. The masses of $B(B^{+},B^{0})$ and $\bar{B}(B^{-},\bar{B^{0}})$ for T=50, 100 and 150 MeV are obtained as 5384.5 (5349.44), 5378.12 (5343.77) and 5374.27 (5340.35) for $\rho_{B}=\rho_{0}$ and 5462.26 (5384), 5453 (5376.87) and 5446.69 (5372) for $\rho_{B}=2\rho_{0}$ in symmetric nuclear matter. For $\rho_{B}=\rho_{0}$, the mass shift of $B$ ($\bar{B}$) in isospin symmetric nuclear matter from the vacuum value of 5277.8 MeV is thus observed to be 106.7 (71.64), 100.32 (66) and 96.47 (62.55), corresponding to $B-\bar{B}$ mass splittings as 35.06, 34.32 and 33.92 respectively. These values are similar to the value of the $D-\bar{D}$ mass splitting of around 32 MeV in isospin symmetric nuclear matter at $\rho_{B}=\rho_{0}$. The mass shifts (in MeV) from the vacuum values for the $B$ and $\bar{B}$ in isospin symmetric nuclear matter ($\sim$ 107 and 71) are observed to be much larger than the values of around 75 and 43 MeV for the $\bar{D}$ and $D$ mesons. The effect of isospin asymmetry is observed to be larger for the antiparticles ($B^{+}$ and $B^{0}$) as compared to the particles ($B^{-}$ and $\bar{B^{0}}$). There is observed to be a drop (rise) in the mass of $B^{+}$ ($\bar{B^{0}}$) for isospin asymmetric (with $\eta$=0.5) nuclear and hyperonic matter as compared to the isospin symmetric cases.
In fig. 11, the Borel curves for the strange-bottom mesons ($B_{s}^{0}$ and $\bar{B_{s}}^{0})$ are shown for density $\rho_{B}=2\rho_{0}$ and Fig. 12 shows the density dependence of these mesons for the nuclear and the strange hadronic matter. The mass of the antiparticle (${B_{s}^{0}}$) is observed to have lower mass as compared to the mass of the particle ($\bar{B_{s}^{0}}$) in nuclear matter ($f_{s}$=0) for which $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$ is zero, whereas, the opposite trend is observed when there is non-zero strangeness in the medium. The lower (higher) mass of the particle ($\bar{B_{s}^{0}}$) as compared to the antiparticle ($B_{s}^{0}$) in the nuclear (hyperonic) matter shows that the term $\langle s^{\dagger}s\rangle$ dominates over the other two terms in the odd part of the spectral function, which is responsible for the splitting of the masses between the particle and antiparticle. The behaviour of the mass splitting at finite densities is similar to the case of strange-charm mesons (as shown in Figs. 5 and 6). Similar to the strange-charm mesons ($D_{s}^{\pm}$), the isospin asymmetry leads to a drop (rise) in the mass of $\bar{B_{s}^{0}}({B_{s}^{0}})$ for nuclear (hyperonic) matter. However, the mass modifications due to isospin asymmetry are observed to be marginal for T=50 and 100 MeV in hyperonic matter.
V Summary
In the present work, we have studied the medium modifications of the masses of the open charm ($D$, $\bar{D}$, $D_{s}$ and $\bar{D}_{s}$) and open bottom ($B$, $\bar{B}$, $B_{s}$ and $\bar{B}_{s}$) in isospin asymmetric strange hadronic matter at finite temperature. These are computed using a QCD sum rule approach retaining the QCD operators upto dimension 5 in the operator product exapnsion. The values of the QCD operators are calculated within a chiral SU(3) model. The mass splitting of the particles ($D^{0}$, $D^{+}$, $D_{s}^{+}$, $B^{-}$, $\bar{B^{0}}$, $\bar{B_{s}^{0}}$) and their antiparticles ($\bar{D^{0}}$, $D^{-}$, $D_{s}^{-}$, $B^{+}$, ${B^{0}}$, ${B_{s}^{0}}$) arises due to the odd part of the spectral function. In nuclear matter, the masses of the particles, $D^{+}(D^{0})$ and $B^{-}(\bar{B^{0}})$ mesons are observed to be smaller than the masses of the antiparticles, $D^{-}({\bar{D}}^{0})$ and $B^{+}({B^{0}})$. For the strange-charm as well as strange-bottom mesons, the opposite behaviour of the masses of particles ($D_{s}^{+}$ and $\bar{B_{s}^{0}}$) to be higher than the masses of antiparticles ($D_{s}^{-}$ and $B_{s}^{0}$) is observed in nuclear matter, whereas, in the strange hadronic matter, the non-zero value of $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$ leads to the masses of the antiparticles ($D_{s}^{-}$ and $B_{s}^{0}$) to be higher than that of $D_{s}^{+}$ and $\bar{B_{s}^{0}}$ in hyperonic medium. The sign of the difference in the masses of the particle and antiparticle ($m_{+}-m_{-}$) is observed to be opposite in nuclear and hyperonic matter, both for the strange-charm as well as strange-bottom mesons, In the mass splitting between the particle and antiparticle arising due to the odd part of the spectral function, the term $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$ for the strange-charm (strange-bottom) mesons thus dominates over the other QCD operators $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$. The effects from density, isospin asymmetry, strangeness as well as temperature on the masses of the open charm (bottom) mesons are observed to be appreciable. These can modify e.g., the production of the charm mesons in asymmetric relativistic heavy ion collision experiments planned at the future facility at GSI, as well as in experiments at J-PARC.